OpenStudy (anonymous):

Let f(x) = (sin(4x))/x and g(x) = 3x2 + 2. Find the limit of f(x) + g(x) as x approaches 0.

OpenStudy (anonymous):

the limit in \(g\) is easy right? replace \(x\) by \(0\)

OpenStudy (anonymous):

yeah, its the limit of f i'm confused on

OpenStudy (anonymous):

so the real question is, what is \[\lim_{x\to 0}\frac{\sin(4x)}{x}\]

OpenStudy (anonymous):

don't you replace it with cosine?

OpenStudy (anonymous):

nah

OpenStudy (anonymous):

gimmick is to multiply top and bottom by \(4\)

OpenStudy (anonymous):

\[4\lim_{x\to 0}\frac{\sin(4x)}{4x}=4\times 1\]

OpenStudy (anonymous):

this assumes you know \(\lim_{x\to 0}\frac{\sin(x)}{x}=1\) which you do

OpenStudy (anonymous):

if you want to waste time you can use l'hopital but it is not necessary here for that matter, you can recognize \[\lim_{x\to 0}\frac{\sin(4x)}{x}\] as the derivative of \(\sin(4x)\) evaluated at 0, i.e. \( 4\cos(0)\)

OpenStudy (whpalmer4):

Or you could directly use L'Hopital and take \[\lim_{x->0} 4\cos(4x)/1 = 4\]

OpenStudy (anonymous):

or you could do what @whpalmer4 said

OpenStudy (anonymous):

@whpalmer4 \to

OpenStudy (anonymous):

so essentially just 4 and then 2 from g?

OpenStudy (anonymous):

\(x\to 0\)

OpenStudy (anonymous):

right

OpenStudy (anonymous):

cool, thanks!

OpenStudy (whpalmer4):

It's all what you remember better...beginners tend to remember the derivative of sin before they remember the limit of sin x/x, people with some EE background remember the sinc function

OpenStudy (anonymous):

true, but often these problems come pre l'hopital, and right after the limit of sine, usually used to prove the derivative to begin with

OpenStudy (whpalmer4):

hmm, as I recall every calc book I ever used did L'Hopital very early on, before tackling trig functions. but no matter, they're in storage, and she isn't using them :-)

OpenStudy (anonymous):

Mathematica calculated 6 as the limit. Refer to the attachment.

OpenStudy (whpalmer4):

yeah, 6 is what we got, too...we decided the limit of the right hand term was 2 right off the bat and concentrated on the left hand term